OpenStudy (anonymous):

What is the 8th term of the geometric sequence –1, 3, –9, …?

OpenStudy (anonymous):

I got 19,683

Parth (parthkohli):

Hmm...see, it's not -3^8, but -3^7. This is because -1 is also included. Got it?

Parth (parthkohli):

3^7 actually

OpenStudy (anonymous):

yeah kind of

Parth (parthkohli):

Because the nth number when n is even is positive.

Parth (parthkohli):

Got it?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow -1 * -3^{n -1} }\) Here I give you the formula ;P

OpenStudy (cwrw238):

common ratio r = 3 / -1 = -3 8th term = a1 * r^7 = -1* (-3)^7

Parth (parthkohli):

3^7 = 2187. 2187 is correct

OpenStudy (anonymous):

oh ok thanks